Megan FACTORED the expression: 35x@- 38-x+8 as (5x-2) (7x-4). But when Jacob applied the FOIL principle and multiple out (5x-2) (7x-4), he got 35x^2 +8; thus, Megan's solution does not appear to check. WHY IS THAT? Please help Megan to understand this better. Explain your reasoning and CORRECTLY factor the orginal expression, if possible. if the expression is prime, so state it.


Now when I worked it out: I got something totally different..
For Megan I got (5x-4) (7x-2) GCF would be equal to 5x-4
Or (5x-2) (7x-4)
5x*7 x-5 x * 4 -2 * 7 x 2 * 4
5x * 7 x- 5 x* 4-2 *2* 7 x 2 * 4
5 * xx x-5 * 4x-14x+8
35 x ^1+1-20x-14x+8
35 x^2-34 x +8
Or have I gotten myself ALL mixed up?????

Answer by Theo(13342)

well,...
you butchered the problem statement so i had to look elsewhere to see what it really was.
the problem is:
megan factored the expression 35x^2 - 38x + 8 and got (5x-2) * (7x-4).
when jacob applied the foil principle, he got 35x^2 - 34x + 8.
what's wrong?
if you use the foil principle to multiply (5x-2) * (7x-4), you will get:
35x^2 - 34x + 8
you are correct there.
if you factor 35x^2 - 38x + 8 yourself, you should find that the factors are:
(5x - 4) * (7x - 2)
when you multiply these factors out, you get:
35x^2 - 38x + 8
this is the correct answer.
megan got the right number to use but she put them in the wrong place.
5x and 7x were correct.
-4 and -2 were correct.
she just placed them wrong.
if you look at the original equation, you will see that:
-2 * -4 = 8
5 * 7 = 35
5*-2 + 7*-4 = -10 + -28 = -38
but........
5*-4 + 7*-2 = -20 + - 14 = -34
bad factors are:
(5x - 2) and (7x - 4)
good factors are:
(5x - 4) and (7x - 2)